Solve the following exponential or logarithmic equation (Show all work)
1.2e^-5x + 2.6 = 3

Answer by vidya pattar(14)

Hello,
1.2 e^(-5x) + 2.6 = 3
first try to get rid of 2.6 from the left side,
subtract 2.6 from the left side ,
1.2 e^(-5x) = 3 - 2.6
1.2 e^(-5x) = 0.4
now divide 1.2 on both sides
e^(-5x) = 0.4 / 1.2
e^(-5x) = 1/3
take log on both sides , we have
-5x = log ( 1/3)
-5x = -0.4771
now divide by -5 on both sides, we get :
x = 0.0954
Thanks & best Regards
Vidya P